OpenStudy (anonymous):

if f(x)=x^2, simplify f(x+h)-f(x)/h

OpenStudy (anonymous):

\[\frac{f(x+h)-f(x)}{h}=\frac{(x+h)^2-x^2}{h}=\frac{x^2+2xh+h^2-x^2}{h}=2x+h\]